# C-DiffSET: Leveraging Latent Diffusion for SAR-to-EO Image Translation with Confidence-Guided Reliable Object Generation

### v2

@@ Appendix B Further Ablation Studies
B.1 Effect of Total Inference Steps
@@
In Fig. 9 and Fig. 10, we justify the selection of $T_{\text{test}}=50$ inference steps in our experiments. While setting the total inference steps to match the training steps ($T=1,000$) ensures high performance, it incurs prohibitive inference times. To balance performance and efficiency, we evaluate the tradeoff between inference time and metric performance by varying the DDIM noise scheduler’s inference steps. For computational feasibility, experiments are conducted on the first 50 samples of the SpaceNet6 dataset. The results indicate significant performance gains for lower step counts ($T_{\text{test}}=1$ to $T_{\text{test}}=50$), but diminishing returns beyond 50 steps, with negligible improvements in metrics such as FID, LPIPS, and PSNR. Based on this analysis, we select total 50 inference steps as the optimal tradeoff, achieving high-quality outputs within practical inference times (3.376 second per 512$\times$512 image).
@@
Figure 11: Qualitative SAR-to-EO image translation on the QXS-SAROPT dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
@@
Figure 12: Qualitative SAR-to-EO image translation on the QXS-SAROPT dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
@@
Figure 13: Qualitative SAR-to-EO image translation on the QXS-SAROPT dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
@@
Figure 14: Qualitative SAR-to-EO image translation on the SAR2Opt dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
@@
Figure 15: Qualitative SAR-to-EO image translation on the SAR2Opt dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
@@
Figure 16: Qualitative SAR-to-EO image translation on the SpaceNet6 dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
@@
Figure 17: Qualitative SAR-to-EO image translation on the SpaceNet6 dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).

### v3

@@ Appendix B Further Ablation Studies
B.1 Impact of Pretrained LDM and C-Diff Loss
We provide in Table 8 further experiments on additional dataset to validate $\mathcal{L}_{\text{C-Diff}}$. Moreover, Fig. 8 presents visual comparisons of SET results under different loss functions, demonstrating the impact of $\mathcal{L}_{\text{C-Diff}}$ on structural consistency and perceptual quality.
B.2 Effect of Text Prompts
We provide additional ablation studies on the text prompt in Table 9. Since our datasets do not include text annotations, we use a generic prompt (fixed for all samples), with Stable Diffusion v2.1’s classifier-free guidance rather than a null prompt $\varnothing$.
B.3 Effect of Total Inference Steps
@@
In Fig. 9 and Fig. 10, we justify the selection of $T_{\text{test}}=50$ inference steps in our experiments. While setting the total inference steps to match the training steps ($T=1,000$) ensures high performance, it incurs prohibitive inference times. To balance performance and efficiency, we evaluate the tradeoff between inference time and metric performance by varying the DDIM noise scheduler’s inference steps. For computational feasibility, experiments are conducted on the first 50 samples of the SpaceNet6 dataset. The results indicate significant performance gains for lower step counts ($T_{\text{test}}=1$ to $T_{\text{test}}=50$), but diminishing returns beyond 50 steps, with negligible improvements in metrics such as FID, LPIPS, and PSNR. Based on this analysis, we select a total of 50 inference steps as the optimal trade-off, achieving high-quality outputs within practical inference times (3.376 seconds per 512$\times$512 image).
Appendix C Additional Experimental Details
Since SAR features were incorporated as additional inputs to the U-Net $\psi$, the weights of the $\psi$’s first convolutional layer were initialized by repeating the original weights across the SAR channels. Additionally, the confidence map was initialized to 1 to ensure stable optimization, starting from the standard $\ell_{2}$ loss, and $\tau=\log 2\pi$ [53] was applied. For the QXS-SAROPT dataset, we used the provided original 256$\times$256-sized patches without modification, with a batch size of 64. For the SAR2Opt and SpaceNet6 datasets, we performed random cropping to 512$\times$512-sized patches and set the batch size to 16. Data augmentation techniques, including random horizontal and vertical flips and random rotations (multiples of 90 degrees), were applied during training to improve generalization. To prevent overfitting and mitigate pretrained weight forgetting, we used a small initial learning rate of $3\times 10^{-5}$. To ensure reproducibility, a random seed of 2,025 was fixed for all experiments. Data was split with 80% used for training and 20% for testing across all experiments.
@@ Appendix C Additional Experimental Details
Figure 12: Visual comparison of SET results on the QXS-SAROPT dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
@@
Figure 13: Visual comparison of SET results on the QXS-SAROPT dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
@@
Figure 14: Visual comparison of SET results on the QXS-SAROPT dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
@@
Figure 15: Visual comparison of SET results on the QXS-SAROPT dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
@@
Figure 16: Visual comparison of SET results on the QXS-SAROPT dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
@@
Figure 17: Visual comparison of SET results on the SAR2Opt dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
@@
Figure 18: Visual comparison of SET results on the SAR2Opt dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
Figure 19: Visual comparison of SET results on the SpaceNet6 dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
Figure 20: Visual comparison of SET results on the SpaceNet6 dataset. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.